OpenStudy (anonymous):

if p(A)=p(B)=.15, p(C)=.4, and p(D)=2p(E). find the probs of D and E

OpenStudy (amistre64):

are we to assume that they all add up to 1?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

if so, then the ration of D to E is 1 to 2 1 - A - B - C gives us what is left to divvy up dividing that by 3 allows us to sort it by a ration of 1 to 2. D=1/3 E = 2/3 does that make sense?

OpenStudy (anonymous):

not really

OpenStudy (amistre64):

if they all add to one we have: A+B+C+D+E = 1 A+B+C+D+2D = 1 A+B+C+3D = 1 3D = 1-A-B-C D = (1-A-B-C)/3

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

D is 1/3 of whats left .. therefore E is 2/3 of whats left

OpenStudy (anonymous):

i just checked the back of the text book and it says D=.20, E=.10

OpenStudy (amistre64):

yes, that would be what we get. i mighta seen the E,D parts a little backwards tho :)

OpenStudy (anonymous):

more explanation pls

OpenStudy (amistre64):

ive really explaned it as much as i can ... i need to know what you are seeing it as.

OpenStudy (amistre64):

all this is is adding parts to equal 1

OpenStudy (amistre64):

they give you the value of most of the parts, and a relationship between the 2 unknowns

OpenStudy (anonymous):

tx alot. i,m so grateful